A comic book villain is holding you at gun point and is making you drink a sample of acid. She gives you a beaker with 100ml of a strong acid with pH=5. She also gives you a beaker of a strong base with a pH=10. You can add as much of the strong base to the strong acid as you want, and you must then drink the solution. You'd be best off trying to make the solution neutral before drinking it. How much of the base should you add?
Answer #1

PH of acid = 5

[ H+ ] of acid = 10-pH = 10-5 M

mol of acid = ( 10-5 mol / litre ) x ( 100 x 10-3 litre ) = 10-6 mol

solution will become neutral when 10-6 mol base is added to it.

PH + POH = 14

POH = 14 - 10 = 4

[ OH- ] in base = 10-4 M

volume of base required = 10-6 mol / ( 10-4 mol / litre ) = 10-2 litre

= 0.01 x 1000 = 10 mL base should be added to make solution neutral